A circle is placed in a square with a side length of 12 cm, as shown below. Find the area of the shaded region. use the value 3.14 for r, and do not round your answer. Be sure to include the correct unit in your answer.


1
Expert's answer
2022-04-21T13:45:31-0400


Step 1.

Area of the square = L2

= 12cm x 12cm = 144 cm2

Step 2.

Area of the circle = πr2

Where r = 12/2 = 6cm

π = 3.14

Area of a circle = 3.14 x 62 = 113.04cm2


Step 3.

Area of the shaded region = Area of the square - Area of the circle

= 144cm2 - 113.04cm2 = 30.96cm2

Answer: S = 30.96cm2